# Cross-Cultural Differences in Mental Health Expressions on Social Media

### v3

@@ Abstract
Culture moderates the way individuals perceive and express mental distress. Current understandings of mental health expressions on social media, however, are predominantly derived from WEIRD (Western, Educated, Industrialized, Rich, and Democratic) contexts. To address this gap, we examine mental health posts on Reddit made by individuals geolocated in India, to identify variations in social media language specific to the Indian context compared to users from the Rest of the World (RoW). Our experiments reveal significant psychosocial variations in emotions (sadness in India vs anxiety in RoW), temporal orientation (present-focused in India vs past-focused in the West), and sociocultural aspects (substance use vs work/achievement). Clinical psychologists practicing in India validated the findings and underlined significant overlap in mental health-related concerns observed in social media posts and in-person sessions. This study demonstrates the potential of social media platforms for identifying cross-cultural differences in mental health struggles (e.g. seeking help in India vs seeking peer support in RoW). Future research should investigate how mental health assessment can be culturally adapted to personalize interventions, ensuring equitable mental health care for individuals from all cultural backgrounds.
@@
**footnotetext: The authors are co-first authors.
@@ 1 Introduction
Over 197 million individuals in India are diagnosed with mental health disorders Sagar et al. (2020), a disproportionate majority of whom do not receive treatment Singh (2018). The treatment gap for mental health disorders goes up to 95% in India, which is the highest across Asian countries and more severe compared to the gap of 78% in the United States (US) Murthy (2017); Naveed et al. (2020). Stigma against mental disorders, limited accessibility to affordable mental healthcare, and a lack of awareness of mental health disorders, cumulatively contribute to diagnostic barriers for mental health care Meshvara (2002); Lahariya (2018); Krendl and Pescosolido (2020).
Figure 1: Study overview to investigate the linguistic expressions and topics specific to Indian users posting in Mental Health (MH) related subreddits (See Appendix A) compared to users across the rest of the world (RoW).
A recent study reported that 64% of Indian participants used social media for mental health communication Patki and Iyer (2022). Social media language can reveal risk for mental disorders, through patterns in word choice and emotional tone Chancellor and De Choudhury (2020); Guntuku et al. (2017); Eichstaedt et al. (2018). For instance, increasing use of the first-person singular pronouns reflects excessive self-focus Spasojević and Alloy (2001); Holtzman et al. (2017), and phrases expressing self-deprecation indicate negative self-evaluation Kovacs and Beck (1978). Language-based AI models for screening mental health risks can facilitate timely, non-intrusive targeted assessments, to enable timely support and treatment, especially in under-resourced contexts such as India, alleviating the challenges associated with administering traditional assessment methods such as PHQ-9  World Health Organization (2021).
Mental disorders manifest differently across cultures Manson (1995), however, how is still an unanswered research question. Rai et al. (2024) demonstrated linguistic differences in Facebook posts of White and Black individuals with depression, raising questions about the generalizability of widely acknowledged cognitive impairments across populations. In parallel, there is growing concern about the AI models performance for predicting mental health risks in persons of color Rai et al. (2024); Aguirre et al. (2021). AI systems lacking awareness of cross-cultural differences in the manifestations of mental disorders may lead to misdiagnosis Bailey et al. (2009); Lewis-Fernandez et al. (2005).
@@
In this paper, we examine how mental health expressions specific to Indian context vary compared to the rest of the world. Previously, De Choudhury et al. (2017) examined linguistic patterns in the online language of individuals with depression but grouped India with other developing nations - for example, South Africa. Another related study on understanding Indian population work by Pendse et al. (2019), examined help-seeking patterns rather than mental health markers, to inform social media platform design. Our study is the first to focus on mental health-related expressions in social media language posted by users in India. Additionally, we collaborate with clinical psychologists in India to interpret linguistic trends and themes obtained from social media posts, which bolsters this computational study.
@@
Understanding how the language of people facing mental distress varies across cultures calls for interpretable features comprehensible to stakeholders such as psychologists and policymakers Aguirre et al. (2021); Loveys et al. (2018); Burkhardt et al. (2022); Rai et al. (2024). Psychosocial word categories (e.g., Linguistic Inquiry and Word Count (LIWC)) and topic modeling (word clusters derived using LDA) are thus preferred approaches to examine variations in language correlated with depression Aguirre et al. (2021); Loveys et al. (2018); Burkhardt et al. (2022); Rai et al. (2024). Using a blend of cross-disciplinary language features and machine learning models, we address the overarching question, if and how the mental health expressions of Indian users on social media are different from the rest of the world by answering the following:
@@
  • How do linguistic expressions in Reddit posts of individuals experiencing mental health challenges in India differ from individuals in the rest of the world?
@@
  • Do the linguistic expressions of depression specific to India differ significantly to reliably distinguish individuals from MH-India compared to those from MH-RoW?
This paper bridges two critical gaps from previous literature. First, the paper identifies mental health expressions specific to India, the world’s most populous country, by mining Reddit threads. Second, it engages with clinical psychologists practicing in India to validate the empirical findings, providing a culturally informed assessment of cross-cultural comparisons of mental health expressions.
@@ 2 Data
Reddit offers a platform for individuals to share their mental health journey and seek support anonymously, thereby making it a rich source to understand the symptomatology and sequelae of mental health De Choudhury and De (2014); Park et al. (2018); Gaur et al. (2018); Lokala et al. (2022). Previously, Reddit posts have been used for identifying shifts to suicidal ideation De Choudhury et al. (2016), depression symptoms Gaur et al. (2018); Liu et al. (2023), and the mental health expressions of immigrants Mittal et al. (2023b), among others. Other more widely used social media platforms such as Mx Takatak, Moj, Tiktok in India could be potential alternatives to collect India-specific mental health communication however, they lack APIs for data collection Oosga .
@@
2.1 Subreddits: Mental Health vs Control
We extracted $3,195,310$ posts and comments from mental health-related subreddits (See Appendix A for list of subreddits) using the PushShift API Baumgartner et al. (2020). The largest portion of users (36.1%) were members of r/depression. The subreddits for the Control group were queried across all remaining subreddits external to the mental health subreddits.
Most posts and comments in the dataset were posted between 2019 and 2020. As a preprocessing step, we removed deleted usernames and null messages. Assuming that users posting in India-specific subreddits are likely to be Indians, we identified India-focused subreddits (See Appendix A) and then grouped together the users who posted in these subreddits. We considered users who posted at least 500 words (excluding comments) to ensure sufficient linguistic richness in users’ language for further analysis. We grouped users into four groups:
  • MH-India (4185 users): “Individuals geolocated in India and posting in Mental Health Subreddits”,
  • MH-RoW (5588 users): “Individuals geolocated outside India (i.e., Rest of World) and posting in Mental Health Subreddits”,
  • Control-India (2622 users): “Individuals geolocated in India and not posting in mental health subreddits” and,
  • Control-RoW (5594 users): “Individuals geolocated outside India and not posting in mental health subreddits”.
The first group (MH-India) is our group of interest; the remaining are controls. We limit the scope of this study to Reddit posts to obtain expressions of experiences with mental health challenges rather than interactions (as in comments) with others’ mental health challenges.
@@ 2.2 Geolocation - India vs RoW
We used the geolocation inference approach Harrigian (2018) as a second layer of verification for user location. The geolocation model is a location estimation model that utilizes word usage, the frequency distribution of subreddit submissions, and the temporal posting habits of each user to determine their location. Specifically, we use the pre-trained GLOBAL inference model***https://github.com/kharrigian/smgeo/tree/master#models to geolocate users in our dataset. We removed any users not geolocated to their group based on subreddit classification. For example, users in MH-India who are not geolocated to India and users in MH-RoW who are geolocated to India were removed. This functioned as a two-step verification to ensure that users in MH-India were from India. Ultimately, 1200 users out of the initial 4185 users were left in the MH-India group, and 930 users out of 2622 were left in the Control-India group. The majority of the users in the RoW group were geolocated to the US, UK, Canada, Germany, Australia, and Sweden (See Fig 2), affirming the dominance of West-centric data on Reddit. There are hardly 100 users from non-Western nations. From now on, we use the terms RoW and Western nations interchangeably.
@@ 2.3 Matching Control groups with users in MH-India
3 Methods
@@ 3.1 Language Features
We extracted a diverse set of interpretable language features widely used in Psychology literature to understand the unique markers of mental health among Indians. This includes n-grams, psychosocial word categories from LIWC, and topics (word clusters derived using LDA).
@@
    LIWC-22 is a closed dictionary comprising 102 word categories based on psycho-social states (e.g., Cognition, Social Processes, Affect, etc.) to measure mental states from the language. These word categories in LIWC are counted for each user, and the count is normalized by the total number of 1-grams for each user, thereby representing each user as a vector of 102 normalized psychosocial categories.
@@
    We used Latent Dirichlet Allocation (LDA) Blei et al. (2003) to extract latent topics in users’ timeline data to capture themes behind mental distress specific to a culture. While recent neural topic modeling methods such as BERTopic Grootendorst (2022) show superior predictive accuracy, LDA has been shown to provide qualitatively robust topics Dixon et al. (2022).
    We generated three sets of topics by setting the number of topics $=[200,500,2000]$. We evaluated the topics’ quality using Topic Uniqueness (TU) Nan et al. (2019). TU represents the number of times a set of keywords is repeated across topics; a higher TU corresponds to a rarely repeated word, indicating that topics are diverse, which is favorable. Additionally, three co-authors independently reviewed the quality of topics. We set the number of topics to 2000 based on the automated (See Table 2) and manual evaluation. Human annotators preferred high topic granularity because it highlights subtle linguistic variations across cultures.
Type #Topics TU
Posts Only 200 0.7675
Posts Only 500 0.7162
Posts Only 2000 0.7694
Table 2: Topic Uniqueness (TU) Scores for different numbers of topics.
@@ 3.2 Correlation Analysis
To measure the association between language and the groups (i.e., MH-India vs MH-RoW), we performed ordinary least squares regressions with the feature sets. We calculated Pearson $r$ to measure the association of each feature to each group in a one-vs-all setting. This is important in supporting an understanding of how each feature set correlates with the presence of mental health challenges in India. p-values were corrected using Benjamini-Hochberg correction for multiple hypothesis testing. 23,344 1-3 grams, 102 word categories for LIWC, 2000 LDA topics, were considered for p-value correction.
@@ 4 Results
4.1 Mental Health Expression: India vs RoW
@@ 4.1.1 N-grams
Sixty-one 1-3 grams were significantly ($p<0.05$) correlated with the MH-India group, and 156 were correlated with the MH-RoW group. Figure 4 illustrates the top 30 1-3 grams arranged in decreasing order of Pearson $r$ for MH-India. Introductory phrases (i am, from india), negation (i don’t know), and help-seeking phrases (suggest, please help) are exclusively correlated with MH-India, reflecting the struggle of users in MH-India for accessing mental help. Academic-related stressors (college, exam, study) are exclusively seen in discussions in MH-India. This is particularly interesting, considering users in both groups were matched for age, yet discussions around student-life challenges are prevalent exclusively in MH-India. Self-referential pronouns such as I and me, negative feelings and symptoms mentions are more frequently seen in MH-RoW (see Fig. A1).
Figure 5: Pearson r for top-20 LIWC categories (with top-3 words) correlated with MH-India with corresponding correlation with MH-RoW. Bars in gray color indicate insignificant correlation ($p>0.05$). p-values were corrected using Benjamini-Hochberg correction. Sadness, social behaviors, and present focus are the most prevalent categories.
Figure 6: Pearson r for top-20 LIWC categories (with top-3 words) correlated with MH-RoW with corresponding correlation with MH-India. Bars in gray color indicate insignificant correlation ($p>0.05$). p-values were corrected using Benjamini-Hochberg correction. Substances, 1st person sing. pronouns, and Mental are the most prevalent categories.
@@ 4.1.2 LIWC
Fifty-two LIWC categories were significantly associated ($p<0.05$) with the MH-India group, whereas sixty categories were found to be correlated with the MH-RoW group. We provide the top LIWC categories for MH-India and MH-RoW in Fig 5 and 6. Sadness, a subset of negative emotions is more strongly correlated with MH-India whereas anxiety (also a type of negative emotion) is more correlated with MH-RoW. Mental health-related discourse is more present-focused in India whereas in MH-RoW group, it is past-focused. This aligns with n-grams findings i.e. MH-India’s emphasis on seeking help and MH-RoW’s on discussing symptoms. Social aspects including the use of 2nd/3rd person pronouns (e.g. you, we), communication, work, politeness, and achievement are positively correlated with only MH-India – reflecting the potential sociocultural expectations (such as collectivism and high-power distance in communication; Robert et al. (2000)) in Indian society.
Discussions in MH-RoW use more clinical language with concepts from psychosocial categories such as mental, health, and illness (also observed in Fig. A1) - reflecting awareness about mental health disorders. Substances are the most strongly correlated psychosocial category with MH-RoW however, it has a negative correlation with MH-India. The correlation with 1st person singular pronouns is also three times more in MH-RoW compared to MH-India. This aligns with recent findings that the use of self-referential pronouns is more likely a marker of depression in European Americans compared to other demographic groups. Similarly, cognitive processes such as all nor none (also a type of cognitive distortion Mercan et al. (2023)) and certitude are more strongly correlated with MH-RoW.
@@ 4.3 Is the language in MH-India different from control groups?
We used a logistic regression model with each feature set (i.e., n-grams, LIWC, and LDA topics) to predict membership in MH-India vs MH-RoW. High AUC scores (See Table 3) demonstrate that users’ language in the MH-India group significantly differs from those in the control groups, including MH-RoW. All language feature groups (i.e., n-grams, LIWC, and LDA topics) have fairly high AUC, indicating differences in mental health expressions at 1-3 gm level, in psycho-social categories as well as in latent topics of discussion.
@@ 5 Discussion
Our study uncovers significant cultural differences in the way users in the MH-India and MH-RoW groups discuss their mental health, underscoring the influence of sociocultural settings on how people perceive mental disorders and seek help. While MH-RoW discussions predominantly focus on feelings, symptoms, and peer support, MH-India situate their struggles with family, education, and work pressures. Although a similar distribution for age and gender is maintained across groups, academic and family pressures are unique to Indian users.
@@
Another contrast observed in Fig 5 is the present-focused language used by Indian social media users, who often express immediate hopelessness or sadness with their current situation Bahri et al. (2023). This differs from the in-person Indian patients, who may ruminate more on past failures or anticipate future challenges. Pendse et al. (2019) also found that Indians discuss “wanting or needing friends” on Mental Health Support Forum more than other countries. We also observe similar behavior in MH-India on Reddit – Indian users are more likely to seek help on social media platforms (see Fig. 4 and 5). Similarly, somatization is the tendency to express emotional distress through bodily symptoms, such as fatigue, headaches, and body pain. This phenomenon is more prevalent in non-Western cultures, due to the heavy emphasis on addressing the tangible physical symptoms compared to the non-tangible mental ones Kirmayer and Young (1998). While this is widely observed among in-person patients in India, we observe a stronger correlation for health-related mentions (such as pain, sick, etc.) with MH-RoW (see Fig. 5 and 7). We speculate that this is likely due to the comorbidity of depression with physical disorders.
Finally, the solution-oriented communication style in MH-India, compared to the situation-descriptive posts in MH-RoW, raises critical questions about the predictive utility of social media language to monitor mental health risks in non-Western contexts. A similar observation is made for Facebook posts where Rai et al. (2024) noted that key markers of depression that is, the use of self-referential language and negative emotions on social media do not reflect depression in Black individuals while it does in White individuals.
Regarding latent themes in Reddit discussions, only $56\%$ of 109 LDA topics correlated with the MH-India group were labeled as prevalent in Indian patients by clinical psychologists. We speculate that some of the topics not labeled as "prevalent" are unseen or emerging themes. Of Top-20 topics in MH-India, "not prevalent" topics revolve around Video Games/Hobbies and Programming/Learning indicating the influence of digital content and growing isolation amongst the undiagnosed young population. These topics could be underrecognized concerns. we also found topics related to Environmental Impact of Energy Sources, Humorous content and reactions, among others. Previous research has suggested that people (particularly young people) are increasingly climate anxious and that humor on social media is often used to cope with mental health challenges Schneider (2018); Sanson (2022). Furthermore, certain pop culture references may be crucial to understanding the narrative - as one clinical psychologist pointed out, the top word "Singh" in one of the topics may correspond to the suicide of late Bollywood star Sushant Singh Rajput, a significant event that potentially catalyzed a range of important conversations surrounding mental health in India Akbar et al. (2020).
@@
The growing treatment gap for mental disorders is a major concern in Indian society. The economic loss from mental health conditions between 2012-2030 is estimated at USD 1.03 trillion***United Nations: https://www.who.int/india/health-topics/mental-health. Automated systems that could diagnose and support mental well-being can potentially alleviate the lack of resources, but they would only be useful when designed considering the cultural sensitivities and norms of society. Based on our findings, culturally-aware mental health intervention techniques should seek to bridge the gap to treatment by addressing hypochondriacal ideas and familial embarrassment in particular. This study establishes significant linguistic variations in the mental health-related language in social media posts by Indians compared to individuals from the rest of the world. It thus becomes imperative to test AI models for cultural invariance before their deployment in clinical settings Aguirre et al. (2021); Rai et al. (2024).

### v4

@@ Abstract
Culture moderates the way individuals perceive and express mental distress. Current understandings of mental health expressions on social media, however, are predominantly derived from WEIRD (Western, Educated, Industrialized, Rich, and Democratic) contexts. To address this gap, we examine mental health posts on Reddit made by individuals geolocated in India, to identify variations in social media language specific to the Indian context compared to users from Western nations. Our experiments reveal significant psychosocial variations in emotions and temporal orientation. This study demonstrates the potential of social media platforms for identifying cross-cultural differences in mental health expressions (e.g. seeking advice in India vs seeking support by Western users). Significant linguistic variations in online mental health-related language emphasize the importance of developing precision-targeted interventions that are culturally appropriate.
@@
Sunny Rai∗†, Khushi Shelat∗†, Devansh R Jain, Kishen Sivabalan, Young Min Cho, Maitreyi Redkar, Samindara Sawant, Lyle H. Ungar & Sharath Chandra Guntuku University of Pennsylvania, Carnegie Mellon University, Indian Institute of Technology, Bombay, Disha Counseling Center {sunnyrai, sharathg}@seas.upenn.edu
**footnotetext: These authors contributed equally to this work
@@ 1 Introduction
Over 197 million individuals in India are diagnosed with mental health disorders Sagar et al. (2020), a disproportionate majority of whom do not receive mental healthcare Singh (2018). Generative AI technologies can facilitate affordable and easily accessible mental health assessment and support, especially in under-resourced contexts such as India Stade et al. (2024).
Mental disorders, however, manifest differently across cultures Manson (1995). For example, AI models trained on Black individuals’ language fail to detect depression in Black individuals Rai et al. (2024). Moreover, the frequency of language markers such as self-focus and negative evaluation indicative of depression do not increase with depression in Black individuals’ language. AI systems lacking awareness of cross-cultural differences in mental health communications may lead to misdiagnosis and health inequity Bailey et al. (2009); Lewis-Fernandez et al. (2005).
@@
In this paper, we examine how mental health expressions specific to Indian context vary compared to the rest of the world (RoW). Previously, De Choudhury et al. (2017) examined depression patterns in the online language of individuals from majority world countries including India. Another related work examined help-seeking patterns to inform social media platform design Pendse et al. (2019). This paper bridges two critical gaps from previous literature. First, we focus on mental health expressions specific to India, the world’s most populous country, by mining Reddit threads. Second, we collaborate with clinical psychologists practicing in India to validate the empirical findings, providing a culturally informed assessment of cross-cultural comparisons of mental health expressions.
@@
To inform the research on culturally competent mental health models Sue (1998), we adopt interpretable features that are comprehensible to stakeholders such as psychologists and policymakers, for modeling cross-cultural variations in mental health language. We use psychosocial word categories (e.g., Linguistic Inquiry and Word Count (LIWC)) and topic modeling (word clusters derived using LDA) to examine variations in language correlated with depression Aguirre et al. (2021); Loveys et al. (2018); Burkhardt et al. (2022); Rai et al. (2024). Using a blend of cross-disciplinary interpretable language features and machine learning models, we address the overarching question, if and how the mental health expressions of Indian users on social media are different from the rest of the world (RoW) by answering the following:
@@
  • How do linguistic expressions in Reddit posts of individuals experiencing mental health challenges in India differ from individuals in the RoW?
@@
  • Do the linguistic expressions of depression specific to India differ significantly to distinguish individuals from India compared to those from RoW?
@@ 2 Data
Reddit offers a platform for individuals to share their mental health journey and seek support anonymously, thereby making it a rich source to understand the symptomatology and sequelae of mental health. India ranked 3rd in Reddit’s website traffic globally after the US and the UK in 2024***https://worldpopulationreview.com/country-rankings/reddit-users-by-country. Previously, Reddit posts have been used for identifying shifts to suicidal ideation De Choudhury et al. (2016), depression symptoms Liu et al. (2023), and the mental health expressions of immigrants Mittal et al. (2023b), among others.
2.1 Subreddits: Mental Health vs Control
We extracted $3,195,310$ posts and comments from mental health-related subreddits (See Appendix A for list of subreddits) using the PushShift API Baumgartner et al. (2020). The largest portion of users (36.1%) were members of r/depression. We queried subreddits external to the mental health subreddits to create a control group. We removed deleted usernames and null messages. We considered users who posted at least 500 words (excluding comments) to ensure sufficient linguistic richness in users’ language for further analysis. We limit the scope of this study to Reddit posts to obtain expressions of experiences with mental health challenges rather than interactions (as in comments) with others’ mental health challenges.
@@ 2.2 Geolocation - India vs RoW
For labeling users’ location, we performed two-level analysis. First, we identified India-focused subreddits (See Appendix A) since users posting in India-specific subreddits are likely to be Indians and grouped together users into four groups:
  • MH-India (4185 users): “Individuals posting in India-specific Subreddits and posting in Mental Health Subreddits”,
  • MH-RoW (5588 users): “Individuals NOT posting in India-specific Subreddits and posting in Mental Health Subreddits”,
  • Control-India (2622 users): “Individuals posting in India-specific Subreddits and NOT posting in mental health subreddits” and,
  • Control-RoW (5594 users): “Individuals NOT posting in India-specific Subreddits and NOT posting in mental health subreddits”.
The first group (MH-India) is our group of interest; the remaining are controls.
Second, we used the geolocation inference approach Harrigian (2018) as a second layer of verification for user location. The geolocation model is a location estimation model that utilizes word usage, the frequency distribution of subreddit submissions, and the temporal posting habits of each user to determine their location. Specifically, we use the pre-trained GLOBAL inference model***https://github.com/kharrigian/smgeo/tree/master#models to geolocate users in our dataset. We removed any users not geolocated to their group based on subreddit classification. For example, users in MH-India who are not geolocated to India and users in MH-RoW who are geolocated to India were removed. This functioned as a two-step verification to ensure that users in MH-India were from India.
Manual Evaluation
@@
Ultimately, 1200 users out of the initial 4185 users were left in the MH-India group, and 930 users out of 2622 were left in the Control-India group. The majority of the users ($\approx 95\%$) in the RoW group were geolocated to Western nations (See Fig 1), affirming the dominance of West-centric data on Reddit. From now on, we use the terms RoW and Western nations interchangeably.
@@ 2.3 Matching Control groups with users in MH-India
3 Modeling Cross-Cultural Variations in Mental Health Language
@@ 3.1 Language Features
We extracted a diverse set of interpretable language features widely used in Psychology literature to understand the unique markers of mental health among Indians.
@@
    LIWC-22 is a closed dictionary comprising 102 psychosocial categories to measure cognitive processes in language. These word categories in LIWC are counted for each user, and the count is normalized by the total number of 1-grams for each user, thereby representing each user as a vector of 102 normalized psychosocial categories.
@@
    We used Latent Dirichlet Allocation (LDA) Blei et al. (2003) to extract latent topics in users’ timeline data to capture themes behind mental distress specific to a culture. We do not use neural models or embeddings due to their skew toward West-centric data and inferences Havaldar et al. (2023).
    We generated three sets of topics by setting the number of topics $=[200,500,2000]$. We evaluated the topics’ quality using Topic Uniqueness (TU) Nan et al. (2019). TU represents the number of times a set of keywords is repeated across topics; a higher TU corresponds to a rarely repeated word, indicating that topics are diverse, which is favorable. Additionally, three co-authors independently reviewed the quality of topics. We set the number of topics to 2000 based on the automated and manual evaluation. Human annotators preferred high topic granularity because it highlights subtle linguistic variations across cultures.
@@ 3.2 Correlation Analysis
To measure the association between language and the groups (i.e., MH-India vs MH-RoW), we performed ordinary least squares regressions with the feature sets. We calculated Pearson $r$ to measure the association of each feature to each group in a one-vs-all setting. p-values were corrected using Benjamini-Hochberg correction for multiple hypothesis testing. 23,344 1-3 grams, 102 word categories for LIWC, 2000 LDA topics, were considered for p-value correction.
@@ 4 Results
4.1 Mental Health Expressions: India vs RoW
@@ 4.1.1 N-grams
Sixty-one 1-3 grams were significantly ($p<0.05$) correlated with the MH-India group, and 156 were correlated with the MH-RoW group. Figure 3 illustrates the top 30 1-3 grams arranged in decreasing order of Pearson $r$ for MH-India. Introductory phrases (i am, from india), negation (don’t, don’t know), and help-seeking phrases (suggest, please help) are exclusively correlated with MH-India, reflecting the struggle of users in MH-India for accessing mental help. Academic-related stressors (college, exam, study) are exclusively seen in discussions in MH-India. This is particularly interesting, considering users in both groups were matched for age, yet discussions around student-life challenges are prevalent exclusively in MH-India. Self-referential pronouns such as I and me, negative feelings and symptoms mentions are more frequently seen in MH-RoW (see Fig. A1). Overall, the discussion in MH-India subreddits is centered around seeking help, whereas symptoms and diagnosis are more commonly discussed in the MH-RoW group.
Figure 4: Pearson r for top 20 LIWC categories (with top-3 words) correlated with MH-India with corresponding correlation with MH-RoW. Bars in gray color indicate insignificant correlation ($p>0.05$). p-values were corrected using Benjamini-Hochberg correction. Sadness, social behaviors, and present focus are the most strongly correlated categories.
Figure 5: Pearson r for top 20 LIWC categories (with top-3 words) correlated with MH-RoW with corresponding correlation with MH-India. Bars in gray color indicate insignificant correlation ($p>0.05$). p-values were corrected using Benjamini-Hochberg correction. Substances, 1st person sing. pronouns, and Mental are the most strongly correlated categories.
@@ 4.1.2 LIWC
Fifty-two LIWC categories were significantly associated ($p<0.05$) with the MH-India group, whereas sixty categories were found to be correlated with the MH-RoW group. We provide the top LIWC categories for MH-India and MH-RoW in Fig 4 and 5. Sadness, a subset of negative emotions is more strongly correlated with MH-India whereas anxiety (also a type of negative emotion) is more correlated with MH-RoW. Mental health-related discourse is more present-focused in India whereas in MH-RoW group, it is past-focused. This aligns with n-grams findings i.e. MH-India’s emphasis on seeking help and MH-RoW’s on discussing symptoms. Social aspects including the use of 2nd/3rd person pronouns (e.g. you, we), communication, work, politeness, and achievement are positively correlated with only MH-India – reflecting the potential sociocultural expectations (such as collectivism and high-power distance in communication Robert et al. (2000)) in Indian society.
Discussions in MH-RoW use more clinical language with concepts from psychosocial categories such as mental, health, and illness (also observed in Fig. A1) - reflecting awareness about mental health disorders. Substances are the most strongly correlated psychosocial category with MH-RoW however, it has a negative correlation with MH-India. The correlation with 1st person singular pronouns is also three times more in MH-RoW compared to MH-India. This aligns with recent findings that the use of self-referential pronouns is more likely a marker of depression in European Americans compared to other demographic groups Rai et al. (2024). Similarly, cognitive processes such as all or none (also a type of cognitive distortion Mercan et al. (2023)) and certitude are more strongly correlated with MH-RoW.
@@ 4.2 Is the language in MH-India different from control groups?
High AUC scores (See Table 2) demonstrate that users’ language in the MH-India group significantly differs from those in the control groups, including MH-RoW. All language feature groups (i.e., n-grams, LIWC, and LDA topics) have fairly high AUC, indicating differences in mental health expressions at 1-3 gm level, in psychosocial categories as well as in latent topics of discussion.
@@ 5 Discussion
Our study uncovers significant cultural differences in the way users in the MH-India and MH-RoW groups discuss their mental health, underscoring the influence of sociocultural settings on how people perceive mental disorders and seek help. While MH-RoW discussions predominantly focus on feelings, symptoms, and peer support, MH-India situates their struggles with family, education, and work pressures, and are more likely to seek help or advice on social media platforms. Pendse et al. (2019) also found that Indians discuss “wanting or needing friends” on Mental Health Support Forum more than other countries.
@@
Indian social media users often express immediate hopelessness or sadness with their current situation Bahri et al. (2023). This differs from the in-person Indian patients having access to healthcare, who may ruminate more on past failures or anticipate future challenges. Similarly, somatization is widely observed among in-person patients in India Kirmayer and Young (1998), we, however, observe a weaker correlation for health-related mentions (such as pain, sick, etc.) with MH-India compared to MH-RoW. Social media data may underrepresent depression symptoms, as Indian users often seek advice rather than explicitly expressing distress. AI models trained on social media risk missing key depressive indicators, especially in cultures where emotional struggles are less directly verbalized. For instance, De Choudhury et al. (2017) demonstrated that Indian and South Africa-based users are less candid in their posts and less likely to exhibit negative emotions in comparison to their Western counterparts.
Regarding latent themes in Reddit discussions, only $56\%$ of 109 LDA topics correlated with the MH-India group were labeled as prevalent in Indian patients by clinical psychologists.
Of Top-20 topics in MH-India which were labeled as "not prevalent" by psychologists comprise Video Games/Hobbies and Programming/Learning. This indicates the influence of digital content and growing isolation amongst the undiagnosed young population. These topics could be underrecognized concerns.
@@
The growing treatment gap for mental disorders is a major concern in Indian society. The economic loss from mental health conditions between 2012-2030 is estimated at USD 1.03 trillion***United Nations: https://www.who.int/india/health-topics/mental-health. Automated systems that could diagnose and support mental well-being can potentially alleviate the lack of resources, but they would only be useful when designed considering the cultural sensitivities and norms of society. Based on our findings, culturally competent mental health intervention techniques should seek to bridge the gap to treatment by addressing hypochondriacal ideas and familial embarrassment in particular. This study establishes significant linguistic variations in the mental health-related language in social media posts by Indians compared to individuals from the rest of the world.